# Optimal Rates for DP-SCO with a Single Epoch and Large Batches

### v1

@@ Abstract
The most common algorithms for differentially private (DP) machine learning (ML) are all based on stochastic gradient descent, for example, DP-SGD . These algorithms achieve DP by treating each gradient as an independent private query. However, this independence can cause us to overpay in privacy loss because we don’t analyze the entire gradient trajectory. In this work, we propose a new DP algorithm, which we call Accelerated-DP-SRGD (DP stochastic recursive gradient descent), that enables us to break this independence and only pay for privacy in the gradient difference, i.e., in the new information at the current step. Our algorithm achieves the optimal DP-stochastic convex optimization (DP-SCO) error (up to polylog factors) using only a single epoch over the dataset, and converges at the Nesterov’s accelerated rate.
Our algorithm can be run in at most $\sqrt{n}$ batch gradient steps with batch size at least $\sqrt{n}$, unlike prior work which required $O(n)$ queries with mostly constant batch sizes. To achieve this, our algorithm combines three key ingredients, a variant of stochastic recursive gradients (SRG), accelerated gradient descent, and correlated noise generation from DP continual counting. Finally, we also show that our algorithm improves over existing SoTA on multi-class logistic regression on MNIST and CIFAR-10.
@@ 1 Introduction
Differentially private stochastic convex optimization (DP-SCO) and its close counterpart, DP empirical risk minimization (DP-ERM) are two of the most well studied problems in private machine learning [11, 8, 27, 7, 2, 10, 55, 4, 35, 29, 18, 40]. Optimal rates are known for nearly all classes of convex losses, e.g., Lipschitz (and smooth) [8, 27], and Lipschitz with strong convexity [29]. However, there is comparatively little work in optimizing the number of gradient calls [12] required to achieve these rates [35, 27, 29].
In this work, we show that it is possible to obtain optimal DP-SCO error (upto polylog $n,1/\delta$ factors) with a single epoch gradient based algorithm for $\ell_{2}$-Lipschitz and smooth losses. More precisely, we show the following:
There exists a gradient based DP algorithm that makes a single pass over the data set (a.k.a. single epoch  [^4] Formally, a gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$. Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.) and achieves optimal DP-SCO error (within polylog factors), with $T=n/B$ batch gradient steps and batch size $B=O(\sqrt{n})$ The algorithm uses exactly two gradients per training example [^5] Here, $n$ is the number of training examples..
Our work significantly improves the results of [27, 55] that required at least $T=\Omega(n)$ batch gradient steps (a.k.a. adaptive gradient oracle calls) with batch size $B=n/T$ to reach the optimal DP-SCO error. Further, unlike [27] the privacy guarantee of our algorithm does not need to assume convexity, which makes it empirically applicable in non-convex settings too without violating DP guarantees. We demonstrate this with an empirical comparison on multi-class logistic regression problems (for which the clipped loss is non-convex [48]), where we show improvements over the strong baseline algorithm of DP-FTRL (DP follow the regularized leader) [38, 23, 15, 14, 13] on standard benchmarks. (See Section 2 for more details on DP-SGD and DP-FTRL .)
@@
Problem definition: There is a distribution of examples $\mathcal{D}$ and a loss function $f(\mathbf{x},d)$ where $\mathbf{x}\in\mathcal{C}$ is a model in the constraint set $\mathcal{C}$ and $d$ is an example. We assume for any $d\in supp(\mathcal{D})$, $f$ is convex, $L$-Lipschitz, and $M$-smooth over $\mathcal{C}$. We receive a sequence of $n$ i.i.d. examples $D=\{d_{1},d_{2},\ldots,d_{n}\}\sim\mathcal{D}$. Using these examples, our goal is to choose a model that minimizes the population loss function $F(\mathbf{x})=\mathbb{E}_{d\sim\mathcal{D}}\left[f(\mathbf{x},d)\right]$. We will operate in the single-pass streaming setting: we must process examples in the order they appear, and each example appears only once in this order. Finally, the model we output should satisfy $(\varepsilon,\delta)$-DP with respect to the examples. For simplicity, we use the zero-out notion of DP [46]: if $P$ is the distribution of models our algorithm outputs given $\{d_{1},d_{2},\ldots,d_{n}\}$ and $Q$ is the distribution given the same sequence, except one example $d_{i}$ instead has $\nabla f(\mathbf{x},d_{i})=0$ for all $\mathbf{x}$, then we require $\max_{S}[P(S)-e^{\varepsilon}Q(S)],\max_{S}[Q(S)-e^{\varepsilon}P(S)]\leq\delta$.
Adaptive oracle complexity (a.k.a. batch gradient steps): We enforce our algorithm to be single epoch (2.1). Furthermore, we quantify the computational complexity of the algorithms in terms of two quantities: a) Number of adaptive gradient computations ($T$) the algorithm has to perform, i.e., the current computation depends on the output of the gradient computations in the previous state [50, 53], and b) number of gradient oracle calls made to each data point overall. Our algorithm has $T=O(\sqrt{n})$, and makes exactly two gradient calls per data point. In comparison, [27, 55] has $T=O(n)$[27] makes one gradient call/data point (but relies on convexity for privacy guarantee), [55] makes two gradient calls/data point (same as ours). In practice, adaptive calls are much more expensive [6, 37], hence we focus on reducing adaptive computation cost as opposed to focusing on the gap between one vs two gradient oracle calls per data point.
Designing private learning algorithms that minimize adaptive computations is a well-established problem in the literature [50, 53, 21]. Even in practice, adaptive gradient evaluations are more expensive than gradient calls on multiple data points at the same state, especially in distributed learning settings like federated learning [6, 37]. To the best of our knowledge, this is the first paper to explicitly optimize for the adaptive oracle complexity for the DP-SCO problem. [53] show that in the non-private setting, the adaptive oracle complexity for optimal SCO guarantee is $T=\Omega(n^{1/4})$ (with a batch size of $B=O(n^{3/4})$). This is achieved by Nesterov’s accelerated SGD. It is worth pointing out that Accelerated-DP-SRGD does not match this non-private bound even for constant dimensions and $\varepsilon\to\infty$. The main reason being that we use stochastic recursive gradients (SRGs) [45, 55] (instead of minibatch gradients) for privacy purposes, and the variance of the SRGs overwhelm the variance due to DP noise at large $\varepsilon$’s. We leave it as an open problem to match the rates in [53] under DP.
Design of Accelerated-DP-SRGD (DP stochastic recursive gradient descent): In terms of design,
@@
Accelerated-DP-SRGD is based on a careful adaptation of Nesterov’s accelerated stochastic gradient descent [44, 9], combined with the ideas of stochastic recursive gradients (SRG) [45, 55], and matrix factorization (MF) based private continual counting mechanisms [24, 19, 23, 42, 26, 33, 34, 5].
@@
Recall, Nesterov’s acceleration can be expressed as the coupling of two gradient descent based algorithms with different learning rates. More formally, the following are the update rules [^6] Here, $\Pi_{\mathcal{C}}(\cdot)$ is the $\ell_{2}$-projection onto to the constraint set, $\{\eta_{t}\}$ and $\beta$ are scaling parameters, and $\{\tau_{t}\}$ are the coupling parameters.: a) $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}\right)$, b) $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}\right)$, and c) $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$. In our implementation of the Nesterov’s method, we will use a DP method to approximate the gradient computed at $\mathbf{x}_{t}$ computed on a minibatch $\mathcal{B}_{t}$ of size $B$: $\nabla_{t}(\mathbf{x}_{t})=\frac{1}{B}\sum\limits_{i\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d_{i})$. We use the idea of stochastic recursive gradients (SRGs) from [45, 55] to first approximate $\nabla_{t}$ as $\eta_{t}\nabla_{t}\approx\Delta_{t}+\Delta_{t-1}+\ldots$, where each $\Delta_{t}=\eta_{t}\nabla_{t}(\mathbf{x}_{t})-\eta_{t-1}\nabla_{t}(\mathbf{x}_{t-1})$. We set $\eta_{t}=t$, $\beta\approx M\cdot\max\left\{\frac{n^{3/2}}{B^{2}},\frac{n}{B}\right\}$, and $\tau_{t}=\frac{\eta_{t}}{\sum\limits_{\tau\leq t}\eta_{\tau}}$. Following are the few crucial differences from [55]: a) We use a truly accelerated method to achieve the improved batch gradient steps, as opposed to tail-averaged online convex optimization methods (which does not include the term $\mathbf{y}_{t}$), b) because of acceleration, we cannot tolerate higher values of $\eta_{t}$. In particular the results of [55] hold with $\eta_{t}=t^{k}$ (for any $k\geq 1$). To bound the sensitivity of $\Delta_{t}$ for privacy, we need to slow down the learning by choosing $\beta\approx M\sqrt{n}$ (corresponding to the iterations $T=n/B\approx\sqrt{n}$), as opposed $\beta=M$ in the non-private case.
With this choice of parameters, we first show that $\Delta_{t}$ has sensitivity $O(1/B)$. Hence, we can use the standard MF based continual counting mechanisms (Section 2.2) to compute DP variants of each $\eta_{t}\nabla_{t}$ with low error. We finally get an SCO error of the form: $\widetilde{O}\left(\frac{B^{2}}{n^{2}}+\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\right)$, with $B$ being the minibatch size. The first two terms match the non-private optimal SCO error from [41].
Comparison to most related prior work [27, 55]: Two of the prior works which are most related to our work are of [27, 55]. Here, we position the ideas in this paper in regard to these works. For $d,\varepsilon=\Theta(1)$ both these algorithms use $T=O(n)$ batch gradient steps, with batch size $\approx O(1)$. Our main algorithm Accelerated-DP-SRGD , uses Nesterov’s accelerated method [44, 9] as the baseline algorithm to improve on the SoTA to $T=B=O(\sqrt{n})$. One natural question is whether the algorithms in the prior work with small modifications can achieve our adaptive oracle complexity.
[27] either relies on privacy amplification by iteration [28] or localization based methods (which requires adding adaptive $\ell_{2}$-regularization). In either case choosing a batch size larger than $O(\max\{\sqrt{d/n},1\})$ in majority of the batch gradient steps, will result in a sub-optimal DP-SCO error. In comparison, for $M=\Theta(1)$ our algorithm does not have any dependence of $d$ on the batch gradient complexity. Furthermore, [27] heavily uses convexity for their privacy analysis. In contrast, Accelerated-DP-SRGD
can be used seamlessly in non-convex settings without compromising on the DP guarantees (via clipping of $\Delta_{t}$).
@@
The algorithm of [55] and their analysis relies on the access to a low-regret online convex optimization algorithm (OCO) [31]. We know that the regret guarantees of OCO methods cannot be improved beyond $\Omega(1/\sqrt{T})$ even under smoothness [31]. Hence, it is not obvious how the online-to-batch framework [55] can be used to achieve the number of batch gradient steps of our algorithm. One might think of replacing the OCO algorithm in [55] with a SGD method tailored to smooth losses [12, Theorem 6.3] to get around the reduced convergence rate of $\Omega(1/\sqrt{T})$. In Section 4.3 we show that unfortunately, because of the variance introduced by the SRGs, in the absence of an accelerated SGD method, the optimal batch size is still $O(1)$.
@@
[17] provides the best known gradient oracle complexity for optimal DP-SCO, for non-smooth convex losses under $\ell_{2}$-geometry. While their algorithm is not single-epoch, their overall gradient complexity (i.e., the number of gradient oracle calls) is $\widetilde{O}\left(\min\left\{n,\frac{n^{2}}{d}\right\}+\min\left\{\frac{nd^{2}}{\varepsilon},n^{4/3}\varepsilon^{1/3}\right\}\right)$. This result is incomparable to our result, since we rely on smoothness in our analysis, but our overall gradient oracle complexity is an improved $\widetilde{O}\left(\sqrt{n}\right)$. Under constant $\varepsilon$, under all regimes of the dimensionality where the DP-SCO guarantee is meaningful i.e., $d\leq n^{2}$, our gradient oracle complexity is superior. We leave the possibility of removing the smoothness assumption in our paper while obtaining the same (or better) gradient oracle complexity as an open problem. It is also worth mentioning that the results in [17] heavily rely on privacy amplification by sampling [36, 11], and convexity of the loss function for privacy. Our work does not rely on either. At the heart of [17], is a new variance bounded stochastic gradient estimator ReSQue. It is an important open question if this estimator can be used to improve the variance properties of our stochastic recursive gradients (SRGs).

### v2

# Optimal Rates for $O(1)$-Smooth DP-SCO with a Single Epoch and Large Batches

@@ Abstract
In this paper we revisit the DP stochastic convex optimization (SCO) problem. For convex smooth losses, it is well-known that the canonical DP-SGD (stochastic gradient descent) achieves the optimal rate of $O\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p\log(1/\delta)}}{\varepsilon n}\right)$ under $(\varepsilon,\delta)$-DP [11], and also well-known that variants of DP-SGD can achieve the optimal rate in a single epoch [23]. However, the batch gradient complexity (i.e., number of adaptive optimization steps), which is important in applications like federated learning, is less well-understood. In particular, all prior work on DP-SCO requires $\Omega(n)$ batch gradient steps, multiple epochs, or convexity for privacy.
We propose an algorithm, Accelerated-DP-SRGD (stochastic recursive gradient descent), which bypasses the limitations of past work: it achieves the optimal rate for DP-SCO (up to polylog factors), in a single epoch using $\sqrt{n}$ batch gradient steps with batch size $\sqrt{n}$, and can be made private for arbitrary (non-convex) losses via clipping. If the global minimizer is in the constraint set, we can further improve this to $n^{1/4}$ batch gradient steps with batch size $n^{3/4}$. To achieve this, our algorithm combines three key ingredients, a variant of stochastic recursive gradients (SRG), accelerated gradient descent, and correlated noise generation from DP continual counting.
@@ 1 Introduction
Differentially private stochastic convex optimization (DP-SCO) and its close counterpart, DP empirical risk minimization (DP-ERM) are two of the most well studied problems in private machine learning [11, 7, 23, 6, 2, 9, 50, 4, 31, 24, 15, 35]. Optimal rates are known for nearly all classes of convex losses, e.g., Lipschitz (and smooth) [7, 23], and Lipschitz with strong convexity [24]. There also have been extensive work in the recent past [31, 23, 24, 26, 51] optimizing for both the number of gradient oracle calls [12] and rounds of adaptive interactions [46, 48] while achieving optimal DP-SCO error.
In this work, we revisit the problem of reducing both the gradient complexity, and the number of adaptive interactions (a.k.a. batch gradient steps) for smooth [^4] Unless mentioned otherwise, we refer to $O(L/\left\|\mathcal{C}\right\|_{2})$ $\ell_{2}$-smoothness. For brevity, we will sometimes refer to this as $O(1)$-smoothness as this gives smoothness parameter $M=O(1)$ if $L,\left\|\mathcal{C}\right\|_{2}=\Theta(1)$, and $M=O(L/\left\|\mathcal{C}\right\|_{2})$ is within a constant of the lower bound $M\geq L/\left\|\mathcal{C}\right\|_{2}$ that holds if $\mathcal{C}$ contains the global minimizer. and convex losses. We show that it is possible to obtain optimal DP-SCO error (upto $\text{polylog}(n,1/\delta)$ factors), in a single epoch algorithm with sublinear batch gradient steps.
@@
Problem definition: There is a distribution of examples $\mathcal{D}$ and a loss $f(\mathbf{x},d)$ where $\mathbf{x}\in\mathcal{C}\subseteq\mathbb{R}^{p}$ is a model in the constraint set $\mathcal{C}$, $\left\|\mathcal{C}\right\|_{2}\leq R$, and $d$ is an example. We assume for any $d\in supp(\mathcal{D})$, $f$ is convex, $L$-Lipschitz, and $M$-smooth over $\mathcal{C}$. We assume for some $c_{M}=O(1)$, we have $M\leq c_{M}L/\left\|\mathcal{C}\right\|_{2}$. We receive a sequence of $n$ i.i.d. examples $D=\{d_{1},d_{2},\ldots,d_{n}\}\sim\mathcal{D}$. Using these examples, our goal is to choose a model in $\mathcal{C}$ that minimizes the population risk function $F(\mathbf{x})=\mathbb{E}_{d\sim\mathcal{D}}\left[f(\mathbf{x},d)\right]$. We will operate in the single-pass streaming setting: we must process examples in the order they appear, and each example appears only once in this order. Finally, the model we output should satisfy $(\varepsilon,\delta)$-DP w.r.t. the examples. For simplicity, we use the zero-out notion of DP [42]: if $P$ is the distribution of models our algorithm outputs given $\{d_{1},d_{2},\ldots,d_{n}\}$ and $Q$ is the distribution given the same sequence, except one example $d_{i}$ instead has $\nabla f(\mathbf{x},d_{i})=0$ for all $\mathbf{x}$, then we require $\max_{S}[P(S)-e^{\varepsilon}Q(S)],\max_{S}[Q(S)-e^{\varepsilon}P(S)]\leq\delta$.
Our result: We show the following:
  1. 1.
    Unconstrained optimization: Assuming the unconstrained population risk minimizer is in the interior of the constraint set (i.e., $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$, and $\mathbf{x}^{\dagger}\in\mathcal{C}$) [^5] In the unconstrained optimization setting, a standard assumption is a known bound $R$ on the norm of $\mathbf{x}^{\dagger}$ and we can set $\mathcal{C}=B(\boldsymbol{0},R)$ to satisfy our assumption. Hence, our assumption can be viewed as implying an unconstrained optimization of the population risk $F(\mathbf{x})$. A similar assumption was made in [49, Theorem 8] while designing accelerated version of dual averaging methods., there exists a gradient based DP algorithm that makes a single pass over the data set (a.k.a. single epoch) and achieves optimal DP-SCO error (within polylog factors) for smooth and convex losses, with $T=\widetilde{\Theta}(n^{1/4})$ batch gradient steps [^6] Here, $n$ is the number of training examples., batch size $B=n/T$, and using exactly two gradients per training example.
  2. 2.
    Constrained optimization: We also show that even without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$our algorithm obtains the optimal DP-SCO bounds with $T=\sqrt{n}$ batch gradient steps, batch size $B=\sqrt{n}$ and in a single epoch.
Our work uniformly improves over all prior [23, 50, 51] and contemporary [26] works under the assumption that $\mathbf{x}^{\dagger}\in\mathcal{C}$. Under smoothness assumption, prior SoTA (in terms of batch gradient complexity) required gradient complexity of $\tilde{\Theta}(n)$, and with batch gradient complexity of $T=\tilde{\Theta}(\sqrt{n})$. Under a single-epoch constraint, to the best of our knowledge no prior work achieves $T=o(n)$. In comparison, our work improves the gradient complexity to $2n$, and batch gradient steps to $T=\widetilde{\Theta}(n^{1/4}$). A contemporary work [26] matches our batch gradient steps ($T=\tilde{\Theta}(n^{1/4})$). However, they have worse gradient complexity of $\tilde{\Theta}(n^{9/8})$.
Our guarantees, even without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$, improves over all prior work (at least by $\text{polylog}(n)$), and is incomparable to the contemporary work of [26]. (See Table 1.)
Works of [23, 14, 26] also study DP-SCO with higher smoothness conditions of e.g. $M=O(\sqrt{n})$, which by Nesterov smoothing [40] corresponds to the non-smooth setting. However, these works require $n^{\Omega(1)}$ passes over the dataset; to the best of our knowledge there is no result in the literature giving optimal rates for DP-SCO in $n^{o(1)}$ epochs in the non-smooth setting (for all settings of the dimension $p$; [14] uses $O(1)$ epochs if $p\leq\varepsilon^{3/2}$). Since we are interested in the single-epoch setting, we hence only focus on $O(1)$-smoothness, and higher smoothness is beyond scope. As an aside, in the $M=O(\sqrt{n})$ regime, [14] achieves worst-case gradient complexity of $\widetilde{\Theta}(n^{4/3})$, as opposed to $\widetilde{\Theta}(n^{11/8})$ by [26].
Gradient complexity and adaptive interaction complexity (a.k.a. batch gradient steps): We measure the computational complexity of our algorithms under standard complexity measures in the convex optimization literature: a) Gradient complexity [12]: It measures the total number of gradient evaluations of the loss function $f(\mathbf{x},\cdot)$, and b) Batch gradient steps (a.k.a. adaptive interaction complexity)  [^7] Henceforth, we will refer to it as batch gradient steps for consistency. [46, 48]: It corresponds to the number of number of adaptive gradient computations ($T$) the algorithm has to perform, i.e., the current computation depends on the output of the gradient computations in the previous state. Understanding the complexity of an algorithm in terms of number of batch gradient steps have significant attention in the recent past [46, 48, 17, 14]. It is especially important in the context of distributed/federated optimization, where adaptive interactions are much more expensive compared to parallel gradient evaluations at a fixed model state. For example, in cross-device federated learning [33], the number of rounds adaptive server/device communication dominate the overall optimization complexity. Furthermore, our algorithm has additional desirable properties, which are crucial in any practical deployment:
  1. (a)
    Single pass/single epoch: A gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$. Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$. Single epoch  [^8] Notice that the definition does not include any restriction on the memory of the algorithm. While stating our results in the single epoch setting, we additionally incorporate the batch size, and the total number gradients evaluated per data point. algorithms are crucial in learning over data streams (where you can process any data point at any single model state), or in cross-device federated learning (FL) where a device holding a specific data point appears only once during the training [10].
    It is worth emphasizing that there is a crucial difference between an algorithm using a single epoch and an algorithm that performs even two epochs over the data; one may be implementation wise infeasible because of the reasons stated above. So, in Table 1 even if the gradient complexity matches asymptotically between two algorithms, still one may be infeasible to implement due to the fact of not being single epoch.
  2. (b)
    Assuming convexity for privacy: In DP-SCO literature (see Table 1), prior works have assumed that the loss function $f(\mathbf{x},d)$ is convex to prove both the privacy and utility guarantee of their algorithms. Assuming convexity for differential privacy guarantees extremely restricts the applicability of the algorithms in practical scenarios where provable convexity property may not hold (e.g., training non-convex models with DP-SGD [1], or simply clipping the gradients to a fixed $\ell_{2}$-norm [45].)
Surprisingly, our algorithms, even being single epoch and not assuming convexity for privacy, uniformly beat the prior SoTA in terms of gradient complexity and batch gradient steps.
Algorithms Gradient Batch gradient Single Convexity
complexity steps epoch for privacy
Ours (assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$) $2n$ $\widetilde{\Theta}(n^{1/4})$ Yes No
Ours (no assumption) $2n$ $\sqrt{n}$ Yes No
[23] $n$ $\widetilde{\Theta}(n)$ Yes Yes
[50] $2n$ $n$ Yes No
[51] $\widetilde{\Theta}(n)$ $\widetilde{\Theta}(\sqrt{n})$ No Yes
[26] $\widetilde{\Theta}(n^{9/8})$ $\widetilde{\Theta}(n^{1/4})$ No No
Table 1: Comparison of gradient complexity and batch gradient steps. Here, $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$ is the unconstrained population risk minimizer, and the Lipschitz constant ($L$) and the smoothness required for all results is $O(1)$. $\widetilde{\Theta}(\cdot)$ hides $\text{polylog}(n)$ factors.
@@
Our algorithm [Accelerated-DP-SRGD (DP stochastic recursive gradient descent)]: In terms of design, our algorithm is based on a careful adaptation of Nesterov’s accelerated stochastic gradient descent [38, 8], combined with the ideas of stochastic recursive gradients (SRG) [41, 50], and private continual counting mechanisms [19, 16, 18, 37, 22, 29, 30, 5]. In the hindsight, Accelerated-DP-SRGD is arguably more natural than prior/contemporary adaptations of accelerated methods [26, 51] towards reducing the computational complexity of obtaining optimal DP-SCO guarantees.
@@
Recall, Nesterov’s acceleration can be expressed as the coupling of two gradient descent based algorithms with different learning rates. More formally, the following are the update rules [^9] Here, $\Pi_{\mathcal{C}}(\cdot)$ is the $\ell_{2}$-projection onto to the constraint set, $\{\eta_{t}\}$ and $\beta$ are scaling parameters, and $\{\tau_{t}\}$ are the coupling parameters.: a) $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}\right)$, b) $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}\right)$, and c) $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$. In our implementation of the Nesterov’s method, we will use a DP method to approximate the gradient computed at $\mathbf{x}_{t}$ computed on a minibatch $\mathcal{B}_{t}$ of size $B$: $\nabla_{t}(\mathbf{x}_{t})=\frac{1}{B}\sum\limits_{i\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d_{i})$. We use the idea of stochastic recursive gradients (SRGs) from [41, 50] to first approximate $\nabla_{t}$ as $\eta_{t}\nabla_{t}\approx\Delta_{t}+\Delta_{t-1}+\ldots$, where each $\Delta_{t}=\eta_{t}\nabla_{t}(\mathbf{x}_{t})-\eta_{t-1}\nabla_{t}(\mathbf{x}_{t-1})$. We set $\eta_{t}=t$, $\beta\approx M$ if we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$ and $\beta\approx M\cdot\max\left\{\frac{n^{3/2}}{B^{2}},\frac{n}{B}\right\}$ otherwise, and $\tau_{t}=\frac{\eta_{t}}{\sum\limits_{\tau\leq t}\eta_{\tau}}$. Following are the few crucial differences from [50]: a) We use a truly accelerated method to achieve the improved batch gradient steps, as opposed to tail-averaged online convex optimization methods (which does not include the term $\mathbf{y}_{t}$), and b) because of acceleration, we cannot tolerate higher values of $\eta_{t}$. In particular the results of [50] hold with $\eta_{t}=t^{k}$ (for any $k\geq 1$). To bound the sensitivity of $\Delta_{t}$ for privacy, we need to either (i) assume $\mathbf{x}^{\dagger}\in\mathcal{C}$, which allows us to show the gradient steps are decreasing over time and hence the gradient difference between $\mathbf{x}_{t}$ and $\mathbf{x}_{t-1}$ is small by smoothness, or (ii) slow down the learning by choosing $\beta\approx M\sqrt{n}$ (corresponding to the iterations $T=n/B\approx\sqrt{n}$), as opposed to $\beta=M$ in the non-private case.
Our analysis: While we argue above that our algorithm is a natural extension of Nesterov’s accelerated SGD, the privacy and the utility analysis of the algorithm is novel, and far from immediate. We give a brief sketch of our analysis below.
At a high-level our analysis requires bounding the $\ell_{2}$-sensitivity of the gradient difference $\Delta_{t}$ to $O(\frac{1}{B\cdot t})$, and then use the binary tree mechanism [19] to compute DP variants of each $\eta_{t}\nabla_{t}$ with low error. We finally get an SCO error of the form: $\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRB}{n}+\frac{LRB\sqrt{p}}{\varepsilon n^{2}}$, with $B$ being the minibatch size. The first two terms match the optimal DP-SCO error. Setting $B\leq\sqrt{n}$ gives that the third and fourth term also are at most the optimal DP-SCO error. The main challenge is bounding the sensitivity of term $\Delta_{t}$ (stated above). Bounding it naively to $O\left(\frac{tM}{B\beta}+L\right)$ (using the analysis of [50]), may result in a sub-optimal choice of the batch size for the final DP-SCO guarantee. Instead we couple the privacy analysis and the utility analysis together.
Recall, $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$ is the unconstrained minimizer of the population risk $F(\mathbf{x})$, and we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$. We observe that $\left\|\nabla F(\mathbf{x})\right\|_{2}^{2}\leq 2M(F(\mathbf{x})-F(\mathbf{x}^{\dagger}))$ for any $\mathbf{x}\in\mathcal{C}$. This allows us to get a tighter control over the norm of the difference in mini-batch gradients: $\frac{1}{B}\cdot\left\|\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t};d)-\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t-1};d)\right\|_{2}$, rather than bounding it bluntly by $\frac{M}{B\beta}$. As the algorithm proceeds and the excess population risk $F(\mathbf{x})-F(\mathbf{x}^{*})$ goes down, $\left\|\frac{1}{B}\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t};d)\right\|_{2}$ becomes smaller. This in-turn can be used to obtain a stronger bound of $O\left(\frac{M}{B\cdot\beta t}\right)$, on the norm of the difference in mini-batch gradients. We use this to reduce the sensitivity of $\Delta_{t}$ to $O\left(\frac{M}{B\beta}+L\right)$. This is formalized via Lemma 5.5.
If we do not assume that $\mathbf{x}^{\dagger}\in\mathcal{C}$, then we proceed with the naive bound on the gradient difference, and compensate by increasing $\beta$, which in turn requires more iterations for the non-private algorithm to converge and hence a smaller batch size of $\Omega(\sqrt{n})$ (as opposed to $B=\widetilde{\Omega}(n^{3/4})$). This still improves on prior work at least by logarithmic factors. More importantly, we still keep the algorithm single epoch and do not require convexity, which is not possible with the comparable prior work of [51].
Comparison to relevant prior work: While DP-SCO has been studied extensively in the literature, we first focus on the works that are most relevant for this paper [23, 50, 26, 51]. As summarized in Table 1, our bounds are uniformly better than prior and contemporary works under the assumption on the unconstrained population risk minimizer being in the interior of $\mathcal{C}$ (i.e., $\mathbf{x}^{\dagger}\in\mathcal{C}$). Without this assumption, our algorithm is still uniformly better than prior works, and incomparable to the contemporary work of [26]. Unlike many of the prior works, our algorithm is single epoch, and does not require convexity assumption to ensure differential privacy.
@@
In terms of techniqiue, our algorithm heavily relies on the idea of coupling DP-continual observation [16, 19, 13, 29] with stochastcic recursive gradients [41] from [50]. However, there are a few crucial differences which allows us to handle much larger batch sizes while still being a single epoch algorithm.  [50] and their analysis relies on the access to a low-regret online convex optimization algorithm (OCO) [28]. We know that the regret guarantees of OCO methods cannot be improved beyond $\Omega(1/\sqrt{T})$ even under smoothness [28]. Hence, it is not obvious how the online-to-batch framework [50] can be used to achieve the number of batch gradient steps of our algorithm. One might think of replacing the OCO algorithm in [50] with a SGD method tailored to smooth losses [12, Theorem 6.3] to get around the reduced convergence rate of $\Omega(1/\sqrt{T})$. In Section 4.3 we show that unfortunately, because of the variance introduced by the SRGs, in the absence of an accelerated SGD method, the optimal batch size is still $O(1)$. The fact that we use Nesterov’s acceleration allows us to handle much larger batch sizes.
@@
[14] provides the best known gradient oracle complexity for optimal DP-SCO, for non-smooth convex losses under $\ell_{2}$-geometry. While their algorithm is not single-epoch, their overall gradient complexity (i.e., the number of gradient oracle calls) is $\widetilde{O}\left(\min\left\{n,\frac{n^{2}}{p}\right\}+\min\left\{\frac{np^{2}}{\varepsilon},n^{4/3}\varepsilon^{1/3}\right\}\right)$. This result is incomparable to our result, since we rely on smoothness in our analysis, but our overall gradient oracle complexity is an improved $\widetilde{O}\left(\sqrt{n}\right)$. Under constant $\varepsilon$, under all regimes of the dimensionality where the DP-SCO guarantee is meaningful i.e., $p\leq n^{2}$, our gradient oracle complexity is superior. We leave the possibility of removing the smoothness assumption in our paper while obtaining the same (or better) gradient oracle complexity as an open problem. It is also worth mentioning that the results in [14] heavily rely on privacy amplification by sampling [32, 11], and convexity of the loss function for privacy. Our work does not rely on either. At the heart of [14], is a new variance bounded stochastic gradient estimator ReSQue. It is an important open question if this estimator can be used to improve the variance properties of our stochastic recursive gradients (SRGs).